OpenStudy (anonymous):

Help three order of operations sentences you can pick ONE out of them all to do :]

OpenStudy (anonymous):

A scientist is studying global warming. She wants to change her average temperature calculations from Celsius to Fahrenheit. The average temperature that she recorded for the Earth last year was 15 degrees Celsius. The formula for changing from Celsius to Fahrenheit is the quantity of the Celsius temperature times nine-fifths plus 32 degrees. According to the scientist’s data, what was the average temperature, in Fahrenheit, of the Earth last year?

OpenStudy (anonymous):

A scientist wants to estimate the volume of Earth and uses a sphere as a model for the planet. According to NASA, Earth’s diameter is 12,756 km. The volume of a sphere is four-thirds of the product of Pi (π) and the radius cubed. Using 3.14 for π and that the radius of a sphere is equal to two times its diameter, estimate the volume of Earth.

OpenStudy (anonymous):

Those two ^ you can decide to help me with either one 

OpenStudy (ranga):

Let C be Celsius and F be Fahrenheit. Then F = 9/5C + 32 Set C = 15 in the above formula and calculate F

OpenStudy (ranga):

Diameter = 12,756 km Radius = Diameter / 2 = 12756/2 = 6378 km Let V be the volume and R be the radius. V = 4/3(pi)(R^3) Put pi = 3.14, R = 6378 in the formula for V above and compute V. The answer should be in so many km^3 (kilometer cube)

OpenStudy (anonymous):

oK im going along wit hyou

OpenStudy (ranga):

What are you getting for F when C = 15?

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

Thats just 9/5 * 15 right? 27

OpenStudy (ranga):

Don't forget to add the 32 in the formula.

OpenStudy (anonymous):

oK soo 27 + 32 = 59 :

OpenStudy (ranga):

Good! That is 59 degrees Fahrenheit. Now what are you getting for V in the second problem? Note R^3 stands for R exponent 3.

OpenStudy (anonymous):

could you slow down that formula for me?

OpenStudy (ranga):

R^3 = R x R x R = 6378 x 6378 x 6378 = ? 4/3 x (pi) = 4/3 x 3.14 = ? Then multiply the above two results to get your answer.

OpenStudy (anonymous):

Thank you so much :] !MUCH APPRECIATED LIF SAVER <3

OpenStudy (ranga):

You are welcome. V = 1.08 x 10^12 km^3 or \[V = 1.08 \times 10^{12} km ^{3}\]